OpenStudy (anonymous):

Please Help Match each set of measures of a and b with the measure of c that gives a volume of 24 cubic inches. 1.a=2in,b=8in A.c=3in 2.a=4in,b=2in B.c=2in 3.a=3in,b=6in C.c=6in 4.a=6in,b=4in D.c=4in

OpenStudy (anonymous):



OpenStudy (anonymous):

What?????????

jhonyy9 (jhonyy9):

4. is right because 6*4*3=72/3=24

jhonyy9 (jhonyy9):

6*4=24 is the area of base

jhonyy9 (jhonyy9):

multiply c =72/3=24

OpenStudy (anonymous):

a is the slant side not the height

OpenStudy (anonymous):

no a is the